# Exploiting Observation Bias to Improve Matrix Completion

### v2

@@ Abstract
We consider a variant of matrix completion where entries are revealed in a biased manner, adopting a model akin to that introduced by Ma and Chen (2019). Instead of treating this observation bias as a disadvantage, as is typically the case, the goal is to exploit the shared information between the bias and the outcome of interest to improve predictions. Towards this, we consider a natural model where the observation pattern and outcome of interest are driven by the same set of underlying latent or unobserved factors. This leads to a two stage matrix completion algorithm: first, recover (distances between) the latent factors by utilizing matrix completion for the fully observed noisy binary matrix corresponding to the observation pattern; second, utilize the recovered latent factors as features and sparsely observed noisy outcomes as labels to perform non-parametric supervised learning. The finite-sample error rates analysis suggests that, ignoring logarithmic factors, this approach is competitive with the corresponding supervised learning parametric rates. This implies the two-stage method has performance that is comparable to having access to the unobserved latent factors through exploiting the shared information between the bias and outcomes. Through empirical evaluation using a real-world dataset, we find that with this two-stage algorithm, the estimates have 30x smaller mean squared error compared to traditional matrix completion methods, suggesting the utility of the model and the method proposed in this work.
@@ 1 Introduction
Matrix Completion. Matrix completion is a core learning task in many modern applications such as clinical trials Hazewinkel et al. (2022), recommender systems Steck (2010), and public policy evaluation Abadie et al. (2010), to name a few. It is concerned with estimating a matrix from partial and noisy observations of its entries. More specifically, the task consists in estimating an $n\times n$ matrix of interest $\mathbf{X}=(x_{ij})_{(i,j)\in[n]\times[n]}$, where only a noisy subset of its entries is observed. Such subset of entries may be characterized by a mask or a so-called observation matrix $\mathbf{A}=(a_{ij})_{i,j\in[n]\times[n]}$ taking values in $\{0,1\}$, i.e., if $a_{ij}=1$, then that a noisy version $y_{ij}$ of $x_{ij}$ is observed, otherwise the entry $(i,j)$ is considered missing.
Missing Completely At Random (MCAR). Due to its relevance, the matrix completion problem has been extensively studied in the recent years, yet most existing work Keshavan et al. (2010, 2009); Candes and Recht (2012); Candès and Tao (2010); Recht (2011); Negahban and Wainwright (2009); Davenport et al. (2014); Chen and Wainwright (2015); Chatterjee (2015); Xu (2017); Abbe et al. (2020) only focused on the so-called missing completely at random (MCAR) model. In this model, all entries are revealed independently with $a_{ij}\sim\mathsf{Ber}(p)$ for some probability parameter $p$. These works propose consistent estimators provided certain conditions on $p$ are satisfied and $\mathbf{X}$ is of low-rank $d\ll n$ with further structural assumptions. Notably, Keshavan et al. (2010) showed that spectral methods with refinement give consistent estimators provided that $p=\Omega(n^{-1}\log(n))$, and Borgs et al. (2022) proposed an iterative approach based on collaborative filtering which is also consistent in a stronger sense if $p=\Omega(n^{-1+\kappa})$ for some $\kappa>0$.
Missing Not At Random (MNAR). This work departs from the MCAR framework and instead concerns the practical and more general setting where entries may be missing not at random (MNAR). In this model, entries are revealed independently with $a_{ij}\sim\mathsf{Ber}(p_{ij})$ where $p_{ij}$ may even be zero. Indeed, in many real world applications, as is the case in clinical trials Steck (2010) or recommender systems Hazewinkel et al. (2022), the pattern of missing entries may carry some bias that may explain the value of the missing entries. For example, consider the case of restaurant recommendations where $\mathbf{X}$ would correspond to a rating matrix. We would not expect a vegan to visit and rate a steakhouse. Similarly, a recommendation algorithm may not recommend steakhouse to a vegan potentially based on their prior interactions.
Indeed, the presence of such bias in observations is an example of the classical confounding. The factors driving the reasons for not observing an outcome are also influencing the outcome itself. In the setting of matrix completion, however, such confounders are unobserved. This is precisely what has driven recent interest in matrix completion in the MNAR data model. We recall some of the recent and relevant prior works on it next.
Relevant Prior Work on MNAR. We briefly recall recent works from Ma and Chen (2019); Agarwal et al. (2023); Bhattacharya and Chatterjee (2022). Detailed relevant discussions on it can be found therein. We note that Agarwal et al. (2023) is closet to the setup considered here and we shall point to differences between them subsequently.
In Ma and Chen (2019), the propensity matrix $\mathbf{P}=\mathbb{E}[\mathbf{A}]$ is assumed to have low nuclear norm and is estimated directly. It is then used to recover the outcome matrix. Therefore, it requires $p_{ij}$ to be uniformly bounded away from 00, while our work does not impose such a restriction. We also refer an interested reader to other works taking similar approaches Yang et al. (2021); Liang et al. (2016); Schnabel et al. (2016); Wang et al. (2020).
In Bhattacharya and Chatterjee (2022), the authors introduce the self-masking MNAR model, where the observation pattern is a scalar function of the outcomes themselves. Their model can be viewed as a special case of the model introduced in this work. Additionally, our method provides sharper error rates compared to their work. We also refer an interested reader to other works taking similar approach Sportisse et al. (2020b, a).
In Agarwal et al. (2023), the authors consider the model closest to this work where the observation pattern and outcomes are determined by the same set of latent factors. The outcomes have low-rank representation with respect to the latent factors while the observation pattern has generic dependence on the latent factors. However, the dependence of the observation pattern is implicitly constrained by the requirement imposed by their Synthetic Nearest Neighbors algorithm, which requires a certain block pattern of observations. The functional form of this observation pattern is absent beyond such a constraint, thus failing to indicate what functional form would actually induce such patterns.
In comparison, as formally described below, we require that observation patterns have low-rank representation with respect to latent factors while outcomes have a Lipschitz function (with low-rank spectrum) dependence on the latent factor. In a sense, this work can be viewed as a way to formalize both the observation pattern and outcome pattern functionally in the model introduced in Agarwal et al. (2023). Indeed, this allows us to compare the performance of our method with a natural lower bound that is suggested by parametric rate of supervised learning. Such was absent in Agarwal et al. (2023) due to lack of formalizing observation pattern structure.
Contributions.
Below, we summarize the key contributions of our work:
@@
    First, we introduce a novel model in the MNAR framework that is richer, more practical, and more expressive than existing ones. We consider that there is shared information between the missingness pattern and outcomes of interest. More specifically, the probability of observation is $p_{ij}\propto n^{-\beta}(1+\mathbf{u}_{i}^{\top}\mathbf{v}_{j})$ for some $\beta>0$ where $\mathbf{u}_{i},\mathbf{v}_{j}$ are $d$-dimensional latent vectors. The outcome of interest $x_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j})$ where $f$ is a Lipschitz function that admits a low-rank representation of order $d$. We refer the reader to section 2 for the precise presentation of our model.
@@
    Second, we propose a novel two-stage algorithm, referred to as Mask Nearest Neighbors (MNN): (i) in the first stage, we build estimates of the distances between the latent vectors, then use these estimates to perform clustering and denoising of the observations; (ii) in the second stage, we use the revealed entries to impute the missing ones. We establish that MNN achieves, under our shared information model, a finite-sample entry-wise performance guarantee of order $\widetilde{O}(n^{-(2-\beta)/(2d)})$ (see Theorem 4.1).
@@
    Third, the prediction error rates achieved by MNN is competitive with parametric rates for corresponding supervised learning methods despite not knowing the latent vectors, and dealing with an unconventional observation model. More precisely, MNN achieves an error rate of order $\widetilde{O}(N^{-1/(2d)})$ which is comparable to the optimal parametric rate for the corresponding supervised learning task known to be of order $\widetilde{O}(N^{-1/(2+2d)})$ where $N$ denotes the number of observed entries (see Lemma 4.1 and discussion in Section 4.2).
@@
    Finally, we establish empirically the performance gains promised by our theoretical results in the real-world setting. Specifically, we utilize more than a million interaction data points of users on the Glance platform – a smart lock-screen that aims to personalize user experience through recommending dynamic lock screens (also called glances). We find that MNN can improve the mean-squared error by 30x compared to a standard matrix completion method, as can be seen in Table 2. We also report empirical performance using a synthetic dataset to discuss nuanced properties of MNN that are not necessarily captured by theoretical results.
Notation.
We recall $\mathcal{S}^{d-1}$ denotes the unit sphere in $\mathbb{R}^{d}$. We write $[n]$ to mean $\{1,\dots,n\}$. For a given parameter $p\in[0,1]$, $\mathsf{Ber}(p)$ denotes the Bernoulli distribution with parameter $p$. For any $a,b\in\mathbb{R}$, we write $a\wedge b=\min(a,b)$ and $a\vee b=\max(a,b)$. For any given vector $\mathbf{x}\in\mathbb{R}^{d}$, we denote its euclidean by $\lVert\mathbf{x}\rVert$. For a given $m\times n$ matrix $\mathbf{M}$, for any subsets $\mathcal{I}\subseteq[n]$, $\mathcal{J}\subseteq[m]$, we denote $\mathbf{M}_{\mathcal{I},\mathcal{J}}$ the sub-matrix of $\mathbf{M}$ constructed by removing the rows $[n]\backslash\mathcal{I}$ and columns $[m]\backslash\mathcal{J}$. We further use $\mathbf{M}_{i,:}$ (resp. $\mathbf{M}_{:,j}$) to denote the $i^{\text{th}}$ row (resp. $j^{\text{th}}$ column) of $\mathbf{M}$. We denote the max norm of $\mathbf{M}$ by $\|\mathbf{M}\|_{\max}=\max_{(i,j)\in[m]\times[n]}|\mathbf{M}_{i,j}|$, its operator norm by $\|\mathbf{M}\|_{\textup{op}}$, its Frobenius norm by $\|\mathbf{M}\|_{\textup{F}}$, and its two-to-infinity norm by $\|\mathbf{M}\|_{2\to\infty}=\max_{i\in[m]}\|\mathbf{M}_{i,:}\|$. The pseudo-inverse of the matrix $\bm{M}$ is denoted by $\bm{M}^{\dagger}$.
2 Problem Description
2.1 Model and learning task
@@
We posit that users (rows) and items (columns) can be fully described by $d$-dimensional unit vectors where $d\ll n$. More specifically, each user $i$ (resp. item $j$) is associated with a unit vector $\mathbf{u}_{i}\in\mathcal{S}^{d-1}$ (resp. $\mathbf{v}_{j}\in\mathcal{S}^{d-1}$), which we assume to be distributed independently and uniformly at random from the unit sphere. We will refer to $(\mathbf{u}_{i})_{i\in[n]}$ (resp. $(\mathbf{v}_{j})_{j\in[n]}$) as the user (resp. item) latent factors. Such latent factors determine both the probability of observation as well as the outcome of interest.
@@
where the mappings $\theta(\cdot)=(\theta_{1}(\cdot),\dots,\theta_{d}(\cdot))$ and $\phi(\cdot)=(\phi_{1}(\cdot),\dots,\phi_{d}(\cdot))$ are defined from $\mathcal{S}^{d-1}$ to $[-B,B]^{d}$, and are assumed to be $L$-lipschitz. As such, the outcomes, despite forming a low-rank matrix of rank $d$, are not simply bi-linear forms of the user and item latent factors as is traditionally the case in matrix completion.
@@
where $\rho_{n}\in(0,1]$ is a sparsity factor that may vary with $n$. This means that the observation pattern induced is very closely related to the random dot product graph construction Athreya et al. (2018). Considering the infamous selection bias issue in recommender systems, which is well-displayed in Ma and Chen (2019), this is a more realistic assumption. The observation indicator is then $a_{ij}\sim\mathsf{Ber}(p_{ij})$ and independent across user-item pairs. Having $a_{ij}=1$ means that the entry $(i,j)$ is revealed, and we observe a noise-corrupted outcome $y_{ij}$. More precisely, we have
@@
where we write $\star$ to mean that an entry is not revealed. $(\xi_{ij})_{i,j\in[n]}$ are i.i.d. copies of a zero-mean and $\sigma^{2}$-sub-gaussian random variable.
@@
Our objective is that of prediction error minimization: Given the observation pattern $(a_{ij})_{i,j\in[n]}$, and observations $\mathcal{D}=\{y_{ij}:a_{ij}=1,i,j\in[n]\}$, we aim at predicting the outcomes $(x_{ij})_{i,j\in[n]}$ with finite-sample error rates at a user-item level. More specifically, we wish to construct estimates $(\hat{x}_{ij})_{i,j\in[n]}$ with small prediction error $|\hat{x}_{ij}-x_{ij}|$, for all $i,j\in[n]$, in high probability.
In the subsequent sections, we will often prefer matrix notations. We denote $\mathbf{U}^{\top}=\begin{bmatrix}\mathbf{u}_{1}&\cdots&\mathbf{u}_{n}\end{bmatrix}$, $\mathbf{V}^{\top}=\begin{bmatrix}\mathbf{v}_{1}&\cdots&\mathbf{v}_{n}\end{bmatrix}$ in $\mathbb{R}^{d\times n}$, $\mathbf{X}=(x_{ij})_{i,j\in[n]}$, $\mathbf{P}=(p_{ij})_{i,j\in[n]}$, $\mathbf{A}=(a_{ij})_{i,j\in[n]}$ and $\mathbf{Y}=(y_{ij})_{i,j\in[n]}$ in $\mathbb{R}^{n\times n}$.
2.2 Rationale behind the proposed model

### v3

@@ Abstract
We consider a variant of matrix completion where entries are revealed in a biased manner. We wish to understand the extent to which such bias can be exploited in improving predictions. Towards that, we propose a natural model where the observation pattern and outcome of interest are driven by the same set of underlying latent (or unobserved) factors. We devise Mask Nearest Neighbor (MNN), a novel two-stage matrix completion algorithm: first, it recovers (distances between) the latent factors by utilizing matrix estimation for the fully observed noisy binary matrix, corresponding to the observation pattern; second, it utilizes the recovered latent factors as features and sparsely observed noisy outcomes as labels to perform non-parametric supervised learning. Our analysis reveals that MNN enjoys entry-wise finite-sample error rates that are competitive with corresponding supervised learning parametric rates. Despite not having access to the latent factors and dealing with biased observations, MNN exhibits such competitive performance via only exploiting the shared information between the bias and outcomes. Finally, through empirical evaluation using a real-world dataset, we find that with MNN, the estimates have 28x smaller mean squared error compared to traditional matrix completion methods, suggesting the utility of the model and method proposed in this work.
@@ 1 Introduction
Matrix completion is a fundamental task in modern machine learning that has garnered widespread interest across many societal and engineering domains [19, 39, 3, 18]. In matrix completion, the goal is to estimate a matrix from partial and noisy observations of its entries. Specifically, we aim to recover an $m\times n$ matrix $\mathbf{X}=(x_{ij})_{(i,j)\in[m]\times[n]}$ with access to noisy observations of only a small subset of its entries. This subset of observed entries can be characterized by a mask or what we refer to as the observation matrix $\mathbf{A}=(a_{ij})_{i,j\in[m]\times[n]}$ taking values in $\{0,1\}$, where $a_{ij}=1$ if a noisy version $y_{ij}$ of $x_{ij}$ is observed, and 00 otherwise (entry $(i,j)$ is missing) for $(i,j)\in[m]\times[n]$.
The MCAR framework.
Much of the existing literature on matrix completion has been focused on the so-called missing completely at random (MCAR) model. In this framework, all entries are revealed independently $a_{ij}\sim\mathrm{Ber}(p)$, i.e. $a_{ij}=1$ with probability $p$ and 00 otherwise for $p\in(0,1]$. Some early prominent works under the MCAR framework have been based on convex optimization and spectral methods [16, 32, 12, 25, 24, 13, 10]. Specifically, the work of [13] proposed a simple estimation procedure known as Universal Singular Value Thresholding (USVT) that achieves the minimax error rate up to a constant factor under the MCAR framework. In another line of work, non-pararmetric methods based on collaborative filtering have been shown to achieve consistent estimation, notably [29], and [9].
The MNAR framework.
Despite this extensive literature, the MCAR model is typically not applicable in practice, as there are unobserved covariates that are influencing the observation pattern as well as outcome of interest. Indeed, in many real world applications, such as public policy evaluation [3], or recommender systems [19], the observation pattern encodes information that can also help estimate the outcome of interest. As an example, consider a movies recommender system where the rows of $\mathbf{X}$ correspond to users, the columns correspond to movies, and an entry $x_{ij}$ corresponds to a user $i$’s preference of movie $j$. In this scenario, we wish to understand how a user would rate a movie they have not seen, allowing us to provide the user with recommendations for new movies they are likely to enjoy. We note, however, that users are less likely to view movies in genres they do not like, meaning that they are less likely to rate these movies. Clearly the observation pattern for such a matrix is not uniform at random as users’ genre preferences affect which movies they give ratings for. Hence, in order to understand the practical conditions under which matrix completion can be performed, we must focus on settings where observations are missing not at random (MNAR) [28]. Furthermore, we note that the presence of such bias in observations is an example of the classical problem of confounding, meaning, the factors driving the reasons for not observing an outcome are also influencing the outcome itself. In the setting of matrix completion, however, these confounding factors are unobserved. This is exactly what has driven the recent interest in the MNAR framework.
The Question Of Interest.
While there is exciting recent progress in the MNAR framework (see Section 1.1 for details), the following question remains unaddressed:
Can bias in observations, due to the presence of latent factors that influence both the observation pattern and the outcomes, aid in estimating outcomes of interest?
1.1 Prior work on the MNAR framework
Many efforts have been made recently towards addressing the challenges posed by matrix completion in the MNAR framework [30, 4, 8, 44, 26, 36, 43, 35, 34, 1]. These diverse contributions bring unique perspectives on the MNAR framework that are reflected in the differing settings they consider which is indicative of the complexity of the framework itself. Among these studies, we further discuss three works, namely those by [30, 8, 4], that are representative of prior works.
Confounding-agnostic MNAR models.
[30] consider an MNAR model where the so-called propensity score matrix $\mathbf{P}=(p_{ij})_{(i,j)\in[m]\times[n]}=\mathbb{E}[\mathbf{A}]$ [^1] For $(i,j)\in[m]\times[n]$, $p_{ij}$ is the probability of observing the entry $(i,j)$ and is denoted here as propensity score in the language of Causal Inference. is assumed to have low nuclear norm and is estimated directly. This estimate is then used to recover the outcome matrix by leveraging the classical technique of inverse propensity weighing (IPW) [21]. Unfortunately, the use of such technique requires the propensity scores $p_{ij}$ to be uniformly bounded away from 00 in order for their methods to be provably consistent in a mean squared error (MSE) sense. Because they follow a similar approach, other works such as [44, 26, 36, 43, 1] suffer from the same limitation. In contrast, it is worth noting that our work evades such limitation.
[8] consider a self-masking MNAR model where the propensity scores are a scalar function of their corresponding outcomes, i.e., $p_{ij}=\omega(x_{ij})$ with $\omega$ being the scalar function a priori unknown. Under such model, the authors propose modified versions of classic spectral-based methods like USVT and establish their consistency in an MSE sense. [35, 34] follow similar modelling choices and also propose spectral-based methods that are also consistent in the same sense. This collection of works consider a model that can be viewed as a special instance of the model considered in the present work as we shall see in subsequent sections.
None of the considered MNAR models in the aforementioned works account for the presence of confounding in matrix completion, i.e., namely the presence of unobserved confounding factors that can influence both the missiness pattern characterized by $\mathbf{P}$ and outcome of interest $\mathbf{X}$. Consequently, the methods these works propose are agnostic to the presence of confounding, thus leading these works to treat the observation bias as a disadvantage that needs mitigating rather than exploiting as we do. Furthermore, we provide much sharper performance guarantees (entry-wise error bound) compared to MSE. It is worth noting that entry-wise guarantees are much harder to obtain even in the MCAR framework and only recently such results have emerged, e.g. see [6, 11, 9] and references therein.
Confounding-aware MNAR models.
In the nascent literature of matrix completion under the MNAR framework, models that account for the presence of confounding are rather scarce. One work that stands out in this scarce literature is that of [4], where they consider an MNAR model with shared information between the observation pattern and outcomes that is determined by a set of unobserved confounding or latent factors. Specifically, the outcomes have a low-rank representation with respect to the latent factors while the observation pattern has generic dependence on the latent factors. They establish that despite generic dependence between observation pattern and latent factors, as long as there is certain “rectangular” pattern that is part of observed entries, their proposed algorithm Synthetic Nearest Neighbours (SNN) can provide entry-wise performance guarantees. In that sense, the goal of [4] has been to show that it is feasible to estimate outcomes of interest even when observation pattern is confounded. However, this work stops short of utilizing the shared information between observation pattern and outcomes to improve the matrix estimation. And this is exactly the primary motivation for this work.
Indeed, devising matrix completion methods that take advantage of such information is an important aspect that has not been addressed. Addressing it is exactly what our work aims to do, thus already distinguishing our contributions from all prior work on the MNAR framework.
1.2 Contributions
We consider that entries are revealed independently, i.e. $a_{ij}\sim\mathrm{Ber}(p_{ij}),i,j\in[n]$ are mutually independent [^2] In this work, for ease of exposition, we use $n\times n$ matrix. As the reader will notice, the methods and analysis naturally extend for $n\times m$ matrix with $m\neq n$., where the observation probabilities or propensity scores, $\mathbf{P}=(p_{i,j})_{(i,j)\in[n]}$ may take values in $[0,1]$ and vary across rows and columns. If $a_{ij}=1$, then $y_{ij}$ is noisy observation of $x_{ij}$ revealed independently. Starting with this framework, our contributions are listed below:
@@ 1.2 Contributions
    A novel confounding-aware MNAR model. To begin with, we propose a novel model within the MNAR framework for matrix completion where we consider that the observation probabilities $\mathbf{P}$ and outcome of interest $\mathbf{X}$ are determined by the same set of latent factors $(\mathbf{u}_{i})_{i\in[n]},(\mathbf{v}_{j})_{j\in[n]}$, which take values in the unit sphere $\mathcal{S}^{d-1}$ for $d>1$. More specifically, we consider for all $i,j\in[n]$, that $p_{ij}=n^{-\beta}(1+\mathbf{u}_{i}^{\top}\mathbf{v}_{j})/2$ for some $\beta>0$, and $x_{ij}=f(\mathbf{u}_{i},\mathbf{v}_{j})$ for some a priori unknown Lipschitz function $f$. We observe noisy outcomes $\mathbf{Y}=(y_{ij})_{i,j\in[n]}$ where $y_{ij}=\mathbb{E}[x_{ij}]$ if $a_{ij}=1$, otherwise $y_{ij}=\star$, to mean that the entry is missing. We refer the reader to Section 2 for a formal model description and the rationale behind it.
@@
    A confounding-aware matrix completion method with entry-wise performance guarantees. Next, we introduce the Mask Nearest Neighbors (MNN) to estimate $\mathbf{X}$ using the mask matrix $\mathbf{A}$ and the noisy observation matrix $\mathbf{Y}$. The algorithm proceeds in two-stages: (i) estimate distances between the latent vectors using $\mathbf{A}$; (ii) it utilizes $\mathbf{Y}$ along with distance estimates between latent factors to apply non-parametric supervised learning to produce estimate of $\mathbf{X}$. We show that MNN enjoys entry-wise error rates that scale as $\tilde{O}(n^{-(2-\beta)/(2d)})$ provided that $\beta<1/4$ (see Theorem 1).
@@
    Comparing with lower bounds in supervised learning. In the traditional supervised learning setup, the goal is to learn the function $f(\cdot,\cdot)$ using $N$ i.i.d. observations sampled as per a distribution over the domain $\mathcal{S}^{d-1}\times\mathcal{S}^{d-1}$. It is well established that under reasonable conditions, to obtain consistent estimation of the function, it requires $\tilde{\Omega}(N^{-1/(2d)})$ (see Lemma 1) samples. In our setup, on average $N=\Theta(n^{2-\beta})$ observations are available. Therefore, without additional structure it must require $\tilde{\Omega}(N^{-1/(2d)})=\tilde{\Omega}(n^{-(2-\beta)/(2d)})$ observations for consistent estimation of $f$ or $\mathbf{X}$. Indeed, such lower bound would assume that we observe the actual values of latent factors. As noted earlier, we are able to obtain consistent estimation of $\mathbf{X}$ with effectively similar number of observations, i.e. $\tilde{O}(n^{-(2-\beta)/(2d)})$ despite not observing latent factors due to the factorization of $f$. Indeed, the precise number of observations required is less than the lower bound and it is not violating it because of the factorization structure of $f$. The purpose of this discussion is to simply contextualise the non-triviality of our results.
@@
    Empirical validation on a real-world dataset. To establish relevance of the model and subsequently resulting algorithm, MNN, we consider a real-world setting. Specifically, we utilize more than a million interaction data points of users on the Glance [^3] Please note that the dataset we use is not publicly available. However, it can be provided by Glance upon request for verification purposes only. For further information about Glance, we refer the reader to their website: https://glance.com/us. platform – a smart lock-screen that aims to personalize user experience through recommending dynamic lock screens (also called glances). We find that MNN can improve the mean-squared error by 28x compared to a standard matrix completion method (see Table 1). We also report empirical performance using a synthetic dataset to discuss nuanced properties of MNN that are not necessarily captured by theoretical results (see Section 6).
1.3 Notation
We recall that $\mathcal{S}^{d-1}$ denotes the unit sphere in $\mathbb{R}^{d}$. We write $[n]$ to mean $\{1,\dots,n\}$. For a given parameter $p\in[0,1]$, $\mathrm{Ber}(p)$ denotes the Bernoulli distribution with parameter $p$. For any $a,b\in\mathbb{R}$, we write $a\wedge b=\min(a,b)$ and $a\vee b=\max(a,b)$. For any given vector $\mathbf{x}\in\mathbb{R}^{d}$, we denote its euclidean by $\lVert\mathbf{x}\rVert$. For a given $m\times n$ matrix $\mathbf{M}$, for any subsets $\mathcal{I}\subseteq[n]$, $\mathcal{J}\subseteq[m]$, we denote $\mathbf{M}_{\mathcal{I},\mathcal{J}}$ the sub-matrix of $\mathbf{M}$ that contains only the entries ×$\mathcal{I}\times\mathcal{J}$. We further use $\mathbf{M}_{i,:}$ (resp. $\mathbf{M}_{:,j}$) to denote the $i^{\text{th}}$ row (resp. $j^{\text{th}}$ column) of $\mathbf{M}$. We denote the max norm of $\mathbf{M}$ by $\|\mathbf{M}\|_{\max}=\max_{(i,j)\in[m]\times[n]}|\mathbf{M}_{i,j}|$, its operator norm by $\|\mathbf{M}\|_{\textup{op}}$, its Frobenius norm by $\|\mathbf{M}\|_{\textup{F}}$, and its two-to-infinity norm by $\|\mathbf{M}\|_{2\to\infty}=\max_{i\in[m]}\|\mathbf{M}_{i,:}\|$. The pseudo-inverse of the matrix $\mathbf{M}$ is denoted by $\mathbf{M}^{\dagger}$.
2 Model Description
In this section, we present a natural MNAR model that takes into account the presence of unobserved confounding. We provide the reasoning behind this being the natural model.
2.1 The MNAR model
We consider a recommender system setting where the number of users and number of items is equal to $n$. Note that we only make this choice for ease of exposition, but otherwise our results easily extend to the case where the set of items and users are not of equal size.
@@ 2.1 The MNAR model
We posit that users (rows) and items (columns) can be fully described by $d$-dimensional unit vectors where $d\ll n$. More specifically, each user $i$ (resp. item $j$) is associated with a unit vector $\mathbf{u}_{i}\in\mathcal{S}^{d-1}$ (resp. $\mathbf{v}_{j}\in\mathcal{S}^{d-1}$), which we assume to be distributed independently and uniformly at random over the unit sphere. We will refer to $(\mathbf{u}_{i})_{i\in[n]}$ (resp. $(\mathbf{v}_{j})_{j\in[n]}$) as the user (resp. item) latent factors. Such latent factors determine both the probability of observation as well as the outcome of interest.
@@
where the mappings $\theta(\cdot)=(\theta_{1}(\cdot),\dots,\theta_{d}(\cdot))$ and $\phi(\cdot)=(\phi_{1}(\cdot),\dots,\phi_{d}(\cdot))$ are defined from $\mathcal{S}^{d-1}$ to $[-B,B]^{d}$ for some $B>0$, and are assumed to be $L$-lipschitz. As such, the outcomes, despite forming a low-rank matrix, are not simply bilinear forms of the user and item latent factors as is traditionally the case in matrix completion.
@@
where $\rho_{n}\in(0,1]$ is a sparsity factor that may vary with $n$. This means that the observation pattern induced is very closely related to the random dot product graph construction [5], and geometric random graphs [17]. This captures the selection bias in recommender systems [30]. Let $a_{ij}\sim\mathrm{Ber}(p_{ij})$, independent across user-item pairs $i,j\in[n]$: $a_{ij}=1$ means that the entry $(i,j)$ is revealed, and we observe a noise-corrupted outcome $y_{ij}$. Specifically,
@@
where $\star$ indicates that an entry is not revealed, and $(\xi_{ij})_{i,j\in[n]}$ are independent zero-mean and $\sigma^{2}$-sub-gaussian random variables.
@@
Given the observation pattern $(a_{ij})_{i,j\in[n]}$, and observations $\mathcal{D}=\{y_{ij}:a_{ij}=1,i,j\in[n]\}$, produce an estimate $\widehat{\mathbf{X}}=(\hat{x}_{ij})_{(i,j)\in[n]}$ of $\mathbf{X}=(x_{ij})_{i,j\in[n]}$ so that the entry-wise prediction error $|\hat{x}_{ij}-x_{ij}|$, for all $i,j\in[n]$ is as small as possible, with high probability (i.e. probability going to $1$ as $n\to\infty$).
Further notation.
In what follows, we will often prefer matrix notation to ease readability. Thus, we denote $\mathbf{U}^{\top}=\begin{bmatrix}\mathbf{u}_{1}&\cdots&\mathbf{u}_{n}\end{bmatrix}$, $\mathbf{V}^{\top}=\begin{bmatrix}\mathbf{v}_{1}&\cdots&\mathbf{v}_{n}\end{bmatrix}$ in $\mathbb{R}^{d\times n}$, $\mathbf{X}=(x_{ij})_{i,j\in[n]}$, $\mathbf{P}=(p_{ij})_{i,j\in[n]}$, $\mathbf{A}=(a_{ij})_{i,j\in[n]}$ and $\mathbf{Y}=(y_{ij})_{i,j\in[n]}$ in $\mathbb{R}^{n\times n}$.
2.2 Why this model makes sense?